You want to test the sensitivity of a new primer set you are working with, so you decide to perform amplification with the following amounts of DNA added to the PCR: 5 ng, 1 ng, 100 pg, and 10pg. If you know you want to add 4ul of DNA to each reaction, to what concentrations do you need to dilute the DNA to if your stock concentration is 67 ng/uL? Show all your work; volumes should not be less than 2uL and no greater than 500uL
Answer #1

Use formula, C1V1 = C2V2

C1 = stock concentration = 67 ng per ul

V1 = stock volume required = ?

C2 = required concentration = 1 ng, 5 ng, 100 pg, 10 pg

V2 = volume of required solution = 33.5 ul

(We need 4 ul, but we will slightly make more. So, we will make 33.5ul).

1 pg = 0.001 ng

1. 5 ng

67 × V1 = 5 × 33.5

V1 = 2.5 ul

2. 1 ng

67 × V1 = 1 × 33.5

V1 = 0.5 ul

3. 100 pg = 0.1 ng

67 × V1 = 0.1 × 33.5

V1 = 0.05 ul

4. 10 pg = 0.01 ng

67 × V1 = 0.01 × 33.5

V1 = 0.005 ul

(If we will make less than 2 ul volume, than huge volume will be wasted. So, it should not be 2 ul. Its incorrect).

Please rate high.